You are a manager at Percolated Fiber, which is considering expanding its operations in synthetic fiber manufacturing. Your boss comes into your office, drops a consultant's report on your desk, and complains, "We owe these consultants S1.5 million for this report, and I am not sure their analysis makes sense. Before we spend the $28.7 million on new equipment needed for this project, look it over and give me your opinion." You open the report and find the following estimates (in millions of dollars) Project Year Earnings Forecast Sales Revenue 10 29.000 29.000 17.400 17.400 1.600 11.600 2.296 2.296 2.870 2.870 6.434 6.434 2.252 2.252 29.000 29.000 17.400 17400 1.600 11.600 2.296 2.870 6.434 2.252 Cost of Goods Sold Gross Profit - General, Sales and Administrative Expenses - Depreciation 2.296 2.870 2 6.434 2.252 Net Operating Income - Income Tax
- Net Income .182 4.182 .182 4.182 All of the estimates in the report seem correct. You note that the consultants used straight-line depreciation for the new equipment that will be purchased today (year 0), which is what the accounting department recommended. They also calculated the depreciation assuming no salvage value for the equipment. The report concludes that because the project will increase earnings by $4.182 million per year for 10 years, the project is worth $41.82 million. You think back to your glory days in finance class and realize there is more work to be done! First, you note that the consultants have not included the fact that the project will require $10.8 million in working capital up front (year 0), which will be fully recovered in year 10. Next, you see they have attributed S2.296 million of selling, general, and administrative expenses to the project, but you know that $1.148 million of this amount is overhead that will be incurred even if the project is not accepted. Finally, you know that accounting eamings are not the right thing to focus on! a. Given the available information, what are the free cash flows in years 0 through 10 that should be used to evaluate the proposed project? b.げthe cost of capital for this project is 15%, what is your estimate of the value of the new project?
Answer #1

a.

$ in Million
Year Sales Cost of Goods Sold Additional General, Sales and Administrative Expenses* Depreciation Net Profits for Tax Tax Amount**
(35%)
Cash flows from proposed project Working Capital Investment New Equipment Purchase Cost Total free cash flows from proposed project
(A) (B) (C) (D) (A-B-C-D)
0         -                   -                                    -                      -                 -                      -                      -   -10.800 -28.700 -39.500
1 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
2 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
3 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
4 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
5 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
6 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
7 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
8 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
9 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798                   -                   -   7.798
10 29.000 -17.400 -1.148 -2.870 7.582 -2.654 7.798 10.800                 -   18.598

*Additional General, Sales and Administrative Expenses relevant for proposed project = $2.296-1.148 = $1.148 million

**Income Tax Rate = 2.252 / 6.434 = 35%

b. Value of new project (NPV) = -39.500 + 7.798 x PVAF(15%, 9 years) + 18.598 x PVF(15%, 10th year)

= -39.500 + 7.798 X 4.772 + 18.598 X 0.247 = $2.306 million